Room: Xavi, Ignacio
Remote: Denis, Geoff, Elisabetta, Stu
Denis: Service account for shifters configuration: problems configuring Kerberos. Follow up to tickets.
Elisabetta: Running tests with FTS. Everything going smoothly. DC: Fermilab working on writing via Rucio. Gas to start filling the tank middle next week.
Geoff: Nothing to report
Stu: CTA: Will keep current tape format? We understand that is not changing at all.
Comment on FNAL: As soon as Steve deletes data the NP04 data fills the freed space.
Xavi: A note on the availability of space that was commented during last meeting: it's not longer true, at least to the scale that we discussed, but there will be space available around September/October.
Xavi: Granted EOS write access as requested. To check if Elisabetta has admin rights in the managing e-group.